A parallel plate air capacitor is made by using two plates 16cm square, spaced 4.7mm apart. It isconnected to a 12.0V battery. (a) What is the capacitance? (b) What is the charge on each plate? (c)What is the electric field between the plates? (d) What is the energy stored in the capacitor? (e) If the battery is disconnected and then the plates are pulled apart to a separation of 9.4mm, what arethe answers to (a), (b),(c) and (d)?


1
Expert's answer
2021-07-12T12:13:26-0400

a)


"C=\frac{(8.85\cdot10^{-12})(0.16)^2}{0.0047}=48.2\cdot10^{-12}F"

b)


"q=48.2\cdot10^{-12}(12)=5.78\cdot10^{-10}C"

c)


"E=\frac{12}{0.0047}=2550\frac{V}{m}"

d)


"W=0.5(48.2\cdot10^{-12})(12)^2=3.47\cdot10^{-9}\ J"